Figure 1: Wasting space to get the numbers right.... waste this page...

Figure 2: An unstructured mesh M for predicting stresses in a cracked plate.

Figure 3: The underlying graph G vertex-partitioned into 64 blocks.

Figure 4: The corresponding 64-vertex minor graph G′.

Figure 5: A placement and routing of G′ on an 8 × 8 torus.